OpenStudy (anonymous):

sketch the graph of the given function. then state the functions domain and range. y=-1.3(6)^x?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@yoloboy1 ?

OpenStudy (anonymous):

idk im sorry

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@Daveqwerty ??

OpenStudy (anonymous):

@Chris911

OpenStudy (anonymous):

What?!?

OpenStudy (anonymous):

do you know how to solve this?

OpenStudy (anonymous):

@*#$ no

OpenStudy (anonymous):

@Chris911 ??

OpenStudy (anonymous):

@UnkleRhaukus ??

OpenStudy (chris911):

refer to this to set it up and help u out http://www.algebra.com/algebra/homework/Rational-functions/Rational-functions.faq.question.346679.html

OpenStudy (chris911):

@whpalmer4

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@jdoe0001 ??

OpenStudy (jdoe0001):

I'd think you could use a graphing calculator maybe ?

OpenStudy (anonymous):

dont have one @jdoe0001

OpenStudy (jdoe0001):

https://desmos.com/calculator

OpenStudy (anonymous):

\[y=-1.3(6)^x\] when x=0,\[y=-1.3(6)^0=-1.3*1=-1.3\] when x=1 \[y=-1.3(6)^1=-7.8\] when x=2, \[y=-1.3(6)^2=-1.3*36=-46.8\] .................................................. \[when ~x \rightarrow \infty,y=-1.3(6)^\infty \rightarrow-\infty \] when x=-1, \[y=-1.3(6)^{-1}=\frac{ -1.3 }{ 6 }=-0.22\] when x=-2, \[y=-1.3(6)^{-2}=\frac{ -1.3 }{36 }=-.036\] ..................................................................... \[when~x \rightarrow-\infty ,y=-1.3(6)^{-\infty}=\frac{ -1.3 }{6^\infty }=0\] draw these points

OpenStudy (anonymous):

domain=R Range=R

OpenStudy (anonymous):

what does the graph look like @surjithayer

OpenStudy (anonymous):

@whpalmer4 ??